find a point z such that the area under the standard normal curve between 0 and z is .4251 and the value of z os positive.
Answer #1

Solution :

P( 0 < z < z) = 0.4251

P(Z < z) - P(Z < 0) = 0.4251

P(Z < z) = P(Z < 0) + 0.4251

P(Z < z) = 0.5 + 0.4251 = 0.9251

P(Z <  1.44) = 0.9251

z = 1.44